What is the width of a rectangle that has a perimeter of 372 feet and a length of 120 feet?

Answer by ewatrrr(24785)

 
Hi,

P = 2*L + 2*W
372ft = 2*120 + 2w
132ft = 2w
w = 66ft
2*120ft + 2*66ft = 240ft + 132ft = 372t